OpenStudy (anonymous):

Simplify the given expression. (−5 + 3i) • (1 − 2i) −5 − 6i 1 + 13i 2 + i 4 − 5i

OpenStudy (ranga):

Use FOIL method to multiply the two factors and simplify.

OpenStudy (anonymous):

can you please show me how. im really confused. @ranga

OpenStudy (ranga):

(−5 + 3i)(1 − 2i) = -5*1 + (-5) *(-2i) + (3i)(1) + (3i)(-2i) = -5 +10i + 3i - 6(-1) = 1 + 13i

OpenStudy (anonymous):

thank you so much @ ranga

OpenStudy (ranga):

you are welcome.